What's the common ratio and next term of this sequence?
333 1/3, 250, 187 1/2

Answer by Boreal(15235)

250/333 1/3 is the same as 750/1000 or 3/4
check with next pair
187.5/250=3/4.
The next term is 187.5*(3/4)=140 5/8
3/4
140 5/8